A finite population consists of 8 elements.

10,10,10,10,10,12,18,40


A. How many samples of size n = 2 can be drawn this population?

B. List all the possible samples and the corresponding means.

C. Construct the sampling distribution of the sample means.


1
Expert's answer
2021-06-23T17:39:57-0400

A.

N=8

n=2

Number of samples "= \frac{N!}{n!(N-n)!} = \frac{8!}{2!(8-2)!}= 28"

B.



C.